# Information Science Principles of Machine Learning: A Causal Chain Meta-Framework Based on Formalized Information Mapping

### v11

@@ 6.2 Experimental Evaluation of the TVD Upper Bound under Noisy Input and Arbitrary Prediction Strategy
Figure 4: Experimental Evaluation and Results of the TVD Upper Bound under Noisy Input and Arbitrary Prediction Strategy. (a) The upper arc-shaped region of the test dataset is shifted upward, and the resulting vacant region is excluded from the test input. The shifted segment, which does not overlap with the original test dataset, is incorporated as noise into the new test dataset. (b) One hundred groups of samples are randomly selected from the noisy test dataset and processed by the model. The RMSE of the predictions are represented by 100 points along the blue line, while the red line indicates the TVD upper bound between the noisy input distribution and the model distribution. (c) In the case of $P_{l}\geq P_{a}$, 100 samples are randomly selected from the noiseless test dataset, transformed into the corresponding noisy data, and processed by the model. The RMSE of each prediction is represented by 100 points along the blue line, with the red line denoting the TVD upper bound between the noiseless input distribution and the model distribution. (d) In the case of $P_{l}<P_{a}$, 100 samples are randomly selected from the noiseless test dataset, transformed into the corresponding noisy data, and processed by the model. The RMSE of each prediction is represented by 100 points along the blue line, with the red line indicating the TVD upper bound between the noiseless input distribution and the model distribution.
@@ 6.3 Ethical Safety Assurance Experiment
Trial Time w/o Time w/ Time Safe Warning Safe Warning
Number Check (s) Check (s) Ratio Samples Samples Rate Rate
@@
In summary, this ethical safety experiment, grounded in Definition 5 and the formal MLT-MF-based framework for ethical safety, illustrates and validates Theorem 2 regarding the feasibility of integrating ethical safety checking mechanisms and achieving the maximal number of safe outputs.
@@ 6.4.2 Visualization-Based Explanation of the Computational Process
Figure 5: Schematic diagram of the computational process for model M after 165 training steps, showing its handling of the input sample $X=[-1.3535,-1.3535]^{T}$ and output $Y\approx[-0.040,0.277]^{T}$.
@@ 6.4.3 Partition-Based Explanation
Figure 6: Illustration of Prediction Polarity and Impact of Training Steps. (a) Polarity distribution of predicted data around the training set region after model M has been trained for 165 steps. The blue square marks the area where the training dataset is located, centered at (1, 1) with a side length of 1. Green dots indicate samples with all-positive predictions; dark green dots indicate samples with all-positive predictions that also belong to the test set; red dots indicate samples with at least one negative prediction that belong to the test set; orange dots indicate samples with at least one negative prediction that do not belong to the test set. (b) Cumulative effect of up to 400 training steps on the prediction polarity at two points: (1, 1) and (-1.3535, -1.3535). The blue and green lines represent the predicted $y_{1}$ and $y_{2}$ values for the sample (1, 1) from the training set; the red and pink lines represent the predicted $y_{1}$ and $y_{2}$ values for the sample (-1.3535, -1.3535) from the test set.
Fig. 6(a) presents an overall view of numerous inputs processed and predicted by model M. This visualization also represents an ontological state of M as it handles a wide range of input information—the real processing and prediction procedure of M is a direct result of this ontological mapping. Thus, according to Definition 4, Fig. 6(a) constitutes a valid explanation for the distribution of prediction polarities by M.
@@ 7 Discussion and Conclusion
This work, grounded in the Objective Information Theory (OIT), develops the MLT-MF framework, employing a formal logical system to integrate the complete causal chain of model construction, training, input processing, output, and online learning. Under a unified semantics, we propose formal definitions of model interpretability and ethical safety, derive two classes of explicit TVD upper bound theorems, and validate the framework through experiments with feedforward neural network models. The core contributions are as follows:
(1) Enabled Mapping for Information Transmission: We characterize the mathematical mapping and the dual physical-causal nature of information transfer from ontological states to carrier states through the concept of “enabled mapping.” This is further extended to chain transmission, mapping each engineering stage of model development to a segment of the information chain, such that model behavior, training, and inference are reduced to logically verifiable compositions of mappings.
(2) Unified Representation of Interpretability and Safety: By adopting higher-order formal logic as a unified representation of “state,” we propose necessary and sufficient conditions for interpretability—linking input, output, and model within the same formal system via logical closure—and construct maximal safe outputs for ethical safety based on hypergraph maximum independent sets.
(3) Integration of Statistical Error, Noise Structure, and Coverage Geometry: We integrate statistical learning error characterization with information noise structure and training coverage geometry, providing a closed-form TVD upper bound, and operationally decomposing it in terms of coverage fitting error, uncovered mass, KL divergence, and noise deletion/addition terms.
Therefore, this study addresses two longstanding gaps in machine learning theory: the lack of unified semantics and causal chain constraints across different stages, and the absence of verifiable mathematical definitions and constructive guarantees for ethical safety and interpretability. By unifying the “existence–transmission–realization” of information within formal logical explanations, we rigorously establish that model interpretability corresponds to information existence, thereby providing a logical-theoretical foundation for translating ethical safety from normative requirements into decidable and constructible guarantees. Using TVD as a fundamental measure of distributional alignment, and establishing correspondences with practical engineering concerns such as training coverage gaps, uncovered-region prediction strategies, and noise deletion/addition, this work bridges the gap between abstract generalization theory and data coverage/system implementation, thus connecting statistical learning, mathematical logic, information science, and systems engineering.
All conclusions and methods presented are amenable to engineering implementation:
  • Interpretability can be realized at multiple levels through unified formal systems, including mathematical formulas, dataflow diagrams, and parameter representations, which is especially important for high-risk industries.
  • Ethical safety is implemented via conflict detection and maximal independent set approximation of output propositions, enabling a deployable process for maximal safe outputs.
[…]
@@
Of course, this work has limitations: the current formal system is most transparent for finite discrete sets, and further refinement is needed for continuous, time-varying, multimodal, and multi-agent scenarios, particularly regarding measurability and dynamic consistency. The maximal independent set for ethical safety is generally NP-hard, requiring the development of complexity-controlled approximation algorithms and alignment with business ethics standards. TVD upper bounds rely on conservative inequalities such as Pinsker’s; in typical scenarios, error estimates may be tightened using related theoretical advances. Physical causality in enabled mappings calls for interdisciplinary integration with various branches of physics, leveraging parameters such as energy consumption, latency, and uncertainty to advance from qualitative causality to quantitative verification. Future work may include large-scale empirical studies to verify the observability of TVD upper bounds and the benefits of strategy optimization, as well as further development of cross-modal and multi-agent information chain composition and conflict resolution methods, thereby enhancing the universality and practical value of MLT-MF in complex systems.
Acknowledgments and Disclosure of Funding
@@ Appendix A Formal Representation of States
For simplicity, $\mathcal{L}$ does not define connectives such as “and” $\land$, “or” $\vee$, “if and only if” $\leftrightarrow$, or the quantifier “exists” $\exists$, and other commonly used logical symbols. However, mathematical logic theory demonstrates that these symbols are entirely equivalent to appropriate combinations of existing symbols and can make many formulas more concise (Hamilton, 1988). Therefore, they are introduced as defined symbols in subsequent discussions.
@@ Appendix B Finite Automaton States
Using the formal expression of states, we can study in depth the special and important object of finite automata (Editorial Committee of Dictionary of Mathematics, 2002).
@@
Proof: Let $M=\left\langle Q,R,U,\delta,\lambda\right\rangle$ be a finite automaton, and $T=\left\{t_{i}\middle|i=1,\ldots,n\right\}$ the (strictly increasing) time points at which $M$ performs input, output, or state‐transition actions. By finite‐automaton theory (Editorial Committee of Dictionary of Mathematics, 2002), $Q=\left\{q_{t_{i}}\middle|i=1,\ldots,n-1\right\}$, $R=\left\{r_{t_{i}}\middle|i=2,\ldots,n\right\}$, and $U=\left\{u_{t_{i}}\middle|i=1,\ldots,n\right\}$ are nonempty finite sets of input symbols, output symbols, and states, respectively;
@@
$\varphi_{\lambda}\left(t_{i}\right)=\varphi_{U}(t_{i})\land\varphi_{I}(t_{i})\rightarrow\exists\lambda(u_{t_{i}},q_{t_{i}})\land{Eq}^{2}(r_{t_{i+1}},\lambda(u_{t_{i}},q_{t_{i}}))\land{IsElementof}^{2}(r_{t_{i+1}},R)\land\varphi_{O}(t_{i+1})$ (B.5)
@@
describes $M$’s inputs, outputs, and state‐transitions over $T$. Here $Cn(X)$ denotes the logical closure of the set of well-formed formulas $X$. This completes the proof.
@@ C.1 A Simple Feedforward Neural Network Model
We choose a feedforward neural network M with an input layer containing two inputs, a hidden layer containing three neurons, and an output layer containing two outputs (Fig. 7). Its simplicity helps explain the essence of MLT-MF in a clear and straightforward manner.
@@ C.2 Symbol Definitions
MLT-MF uses a formal system to express the operational logic of each phase throughout the model’s full lifecycle. Defining the symbols is a prerequisite for formal modeling. To this end, Tables 2–7 provide definitions for the constants, variables, functions, and predicates used in the formal representation of model M.
@@
Table 3: Definition of Variables - Part I: Logical Formulas and Basic Variables
@@
$\varphi o_{i}$ WFFs well-formed formulas in initial model ontological state
$\psi o_{i}$ WFFs well-formed formulas in initial model carrier state
$\varphi t_{i}$ WFFs well-formed formulas in training ontological state
$\psi t_{i}$ WFFs well-formed formulas in training carrier state
$\varphi u_{i}$ WFFs well-formed formulas in optimized model ontological state
$\psi u_{i}$ WFFs well-formed formulas in optimized model carrier state
$\varphi q_{i}$ WFFs well-formed formulas in input ontological state
$\psi q_{i}$ WFFs well-formed formulas in input carrier state
$\varphi r_{i}$ WFFs well-formed formulas in output ontological state
$\psi r_{i}$ WFFs well-formed formulas in output carrier state
$\varphi v_{i}$ WFFs well-formed formulas in re-optimized model ontological state
$\psi v_{i}$ WFFs well-formed formulas in re-optimized model carrier state
layerslayers natural number number of layers in the model
@@
Table 4: Definition of Variables - Part II: Neural Network Parameters
Symbol Type Definition
$W_{k}^{(o)}$ matrix weight matrix from k-th to (k+1)-th layer (initial)
$w_{ij}^{(o)k}$ real number weight from j-th to i-th neuron (initial)
$b_{k}^{(o)}$ vector bias vector of k-th layer (initial)
$b_{i}^{(o)k}$ real number bias of i-th neuron in k-th layer (initial)
$W_{k}^{(u)}$ matrix weight matrix from k-th to (k+1)-th layer (optimized)
$w_{ij}^{(u)k}$ real number weight from j-th to i-th neuron (optimized)
$b_{k}^{(u)}$ vector bias vector of k-th layer (optimized)
$b_{i}^{(u)k}$ real number bias of i-th neuron in k-th layer (optimized)
$W_{k}^{(v)}$ matrix weight matrix from k-th to (k+1)-th layer (re-optimized)
$w_{ij}^{(v)k}$ real number weight from j-th to i-th neuron (re-optimized)
$b_{k}^{(v)}$ vector bias vector of k-th layer (re-optimized)
$b_{i}^{(v)k}$ real number bias of i-th neuron in k-th layer (re-optimized)
$W_{k}^{(w)}$ matrix weight matrix from k-th to (k+1)-th layer (further re-optimized)
$w_{ij}^{(w)k}$ real number weight from j-th to i-th neuron (further re-optimized)
$b_{k}^{(w)}$ vector bias vector of k-th layer (further re-optimized)
$b_{i}^{(w)k}$ real number bias of i-th neuron in k-th layer (further re-optimized)
Table 5: Definition of Variables - Part III: Vectors and Optimization Parameters
Symbol Type Definition
$X$ vector input vector
$H$ vector hidden layer vector
$Y$ vector output vector
$Y_{t}$ vector output reference vector
$x_{i}$ real number i-th element of input vector
$h_{i}$ real number i-th element of hidden layer vector
$y_{i}$ real number i-th element of output vector
$y_{ti}$ real number i-th element of output reference vector
@@
$\nabla_{W_{1}}L$ real number gradient of loss function w.r.t. $W_{1}$
$\nabla_{b_{1}}L$ real number gradient of loss function w.r.t. $b_{1}$
$\nabla_{W_{2}}L$ real number gradient of loss function w.r.t. $W_{2}$
$\nabla_{b_{2}}L$ real number gradient of loss function w.r.t. $b_{2}$
$\eta$ positive real learning rate
@@
$ReLU(x)$ 1 rectified linear unit (ReLU) activation function
$Identity(x)$ 1 identity function
@@
$MSE(Y,Y_{t})$ 2 mean squared error between vector $Y$ and $Y_{t}$
${ReLU}^{\prime}(x)$ 1 derivative of $ReLU(x)$
$Target(X)$ 1 objective function for input vector $X$
@@
${NetworkStructure}^{2}(layers,k)$ model structure has k layers
${LayerSize}^{2}(k,j)$ the k-th layer has j neurons
${Eq}^{2}(x,y)$ $x$ equals $y$
${InputAt}^{2}(X,t)$ input $X$ at time $t$
${HiddenAt}^{2}(H,t)$ hidden layer value is $H$ at time $t$
${OutputAt}^{2}(Y,t)$ output $Y$ at time $t$
${InputAt}^{3}(X,Y_{t},t)$ input vector $X$ and label vector $Y_{t}$ at time $t$
${Differentiable}^{2}(f,v)$ function $f$ is differentiable with respect to $v$
${BelongsTo}^{2}(x,X)$ element $x$ belongs to set $X$
@@ C.3 Formal Representations of Information at Each Stage of Model M
$Io=\left\langle oo,{To}_{h},S_{oo},M,{To}_{m},S_{M}\right\rangle$ (C.13)
Here, the ontology $oo$ represents the designer’s conceptual blueprint; ${To}_{h}$ is the time at which the model is designed; and the ontological state $S_{oo}(oo,{To}_{h})$ is a set of well-formed formulas that describe the initial operational logic of model M:
@@
$\varphi o_{1}={NetworkStructure}^{2}(layers,3)\land{LayerSize}^{2}(1,2)\land{LayerSize}^{2}(2,3)\land{LayerSize}^{2}(3,2)$ (C.15)
@@
$\varphi o_{2}=\forall i{\forall j(BelongsTo}^{2}\left(i,\left\{1,2,3\right\}\right){\land BelongsTo}^{2}\left(j,\left\{1,2\right\}\right)\rightarrow{\exists v(Eq}^{2}\left(w_{ij}^{(o)1},v\right))$ (C.16)
$\varphi o_{3}=\forall i{\forall j(BelongsTo}^{2}\left(i,\left\{1,2\right\}\right){\land BelongsTo}^{2}\left(j,\left\{1,2,3\right\}\right)\rightarrow{\exists v(Eq}^{2}\left(w_{ij}^{(o)2},v\right))$ (C.17)
@@
$\begin{split}\varphi o_{7}=&\forall X\forall t({InputAt}^{2}(X,t)\land{BelongsTo}^{2}\left(X,\mathbb{R}^{2}\right)\rightarrow\exists H({BelongsTo}^{2}\left(H,\mathbb{R}^{3}\right)\\ &\land{HiddenAt}^{2}(H,t+\delta)\land\forall i({BelongsTo}^{2}\left(i,\left\{1,2,3\right\}\right)\\ &\rightarrow{Eq}^{2}(h_{i},ReLU\left(\sum_{j=1}^{2}w_{ij}^{(o)1}\times x_{j}+b_{i}^{(o)1}\right)))))\end{split}$ (C.21)
$\begin{split}\varphi o_{8}=&\forall H\forall t({HiddenAt}^{2}(H,t+\delta)\land{BelongsTo}^{2}\left(H,\mathbb{R}^{3}\right)\rightarrow\exists Y({BelongsTo}^{2}\left(Y,\mathbb{R}^{2}\right)\\ &\land{OutputAt}^{2}(Y,t+2\delta)\land\forall i({BelongsTo}^{2}\left(i,\left\{1,2\right\}\right)\\ &\rightarrow{Eq}^{2}\left(y_{i},Identity\left(\sum_{j=1}^{3}w_{ij}^{(o)2}\times h_{j}+b_{i}^{(o)2}\right)\right))))\end{split}$ (C.22)
@@
$\begin{split}\varphi o_{9}=&\forall X\forall Y_{t}\forall t({InputAt}^{3}\left(X,Y_{t},t\right)\land{BelongsTo}^{2}\left(Y_{t},\mathbb{R}^{2}\right)\land\varphi o_{7}\land\varphi o_{8}\\ &{\land Eq}^{2}\left(L,MSE\left(Y,Y_{t}\right)\right)\land{Eq}^{2}\left(MSE\left(Y,Y_{t}\right),\frac{1}{2}\sum_{i=1}^{2}\left(y_{i}-y_{ti}\right)^{2}\right))\end{split}$ (C.23)
@@
$\quad\land{Differentiable}^{2}(MSE(Y,Y_{t}),W_{1}^{(o)})$
$\quad\land{Differentiable}^{2}(MSE(Y,Y_{t}),b_{1}^{(o)})$
$\quad\land{Differentiable}^{2}(MSE(Y,Y_{t}),W_{2}^{(o)})$
$\quad\land{Differentiable}^{2}(MSE(Y,Y_{t}),b_{2}^{(o)})$
$\quad\land{Eq}^{2}(\nabla_{W_{1}^{(o)}}L,$
$\qquad({(W_{2}^{(o)})}^{T}(Y-Y_{t})\bigodot{ReLU}^{\prime}(W_{1}^{(o)}X+b_{1}^{(o)}))X^{T})$
$\quad\land{Eq}^{2}(\nabla_{b_{1}^{(o)}}L,$
$\qquad{(W_{2}^{(o)})}^{T}(Y-Y_{t})\bigodot{ReLU}^{\prime}(W_{1}^{(o)}X+b_{1}^{(o)}))$
$\quad\land{Eq}^{2}(\nabla_{W_{2}^{(o)}}L,(Y-Y_{t}){(ReLU(W_{1}^{(o)}X+b_{1}^{(o)}))}^{T})$
$\quad\land{Eq}^{2}(\nabla_{b_{2}^{(o)}}L,(Y-Y_{t}))$ (C.24)
@@
$sup({To}_{h})\leq inf({To}_{m})$ (C.26)
The state $S_{M}(M,{To}_{m})$ is a one-to-one enabling mapping of the state $S_{oo}\left(oo,{To}_{h}\right)$ onto M, that is,

### v12

@@ 6.2 Experimental Evaluation of the TVD Upper Bound under Noisy Input and Arbitrary Prediction Strategy
Figure 4: Experimental Evaluation and Results of the TVD Upper Bound under Noisy Input and Arbitrary Prediction Strategy.
@@ 6.3 Ethical Safety Assurance Experiment
Trial Number
Computation Time
Without Ethical
Safety Check
Computation Time
With Ethical
Safety Check
Computation
Time Ratio
Number of
Ethically Safe
Output Samples
Number of
Unethical
Warning Samples
Correct Prediction
Rate for Safe
Outputs
Correct Prompt
Rate for
Unethical Outputs
@@ 6.4.2 Visualization-Based Explanation of the Computational Process
Figure 5: Schematic diagram of the computational process for model M after 165 training steps.
@@ 6.4.3 Partition-Based Explanation
Figure 6: Illustration of Prediction Polarity and Impact of Training Steps.
@@ 7 Conclusion
This work, grounded in the Objective Information Theory (OIT), develops the MLT-MF framework, employing a formal logical system to integrate the complete causal chain of model construction, training, input processing, output, and online learning. Under a unified semantics, the author proposes formal definitions of model interpretability and ethical safety, derive two classes of explicit TVD upper bound theorems, and validate the framework through experiments with feedforward neural network models.
@@
Of course, this study has limitations: the current formal system is most transparent for finite discrete sets, and further refinement is needed for continuous, time-varying, multimodal, and multi-agent scenarios, particularly regarding measurability and dynamic consistency. The maximal independent set for ethical safety is generally NP-hard, requiring the development of complexity-controlled approximation algorithms and alignment with business ethics standards. TVD upper bounds rely on conservative inequalities such as Pinsker’s; in typical scenarios, error estimates may be tightened using related theoretical advances. Physical causality in enabled mappings calls for interdisciplinary integration with various branches of physics, leveraging parameters such as energy consumption, latency, and uncertainty to advance from qualitative causality to quantitative verification. Future work may include large-scale empirical studies to verify the observability of TVD upper bounds and the benefits of strategy optimization, as well as further development of cross-modal and multi-agent information chain composition and conflict resolution methods, thereby enhancing the universality and practical value of MLT-MF in complex systems.
Acknowledgement
@@
Declaration of competing interest
The author declare that he has no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
@@ Appendix A Formal Representation of States
For simplicity, $\mathcal{L}$ does not define connectives such as “and” $\land$, “or” $\vee$, “if and only if” $\leftrightarrow$, or the quantifier “exists” $\exists$, and other commonly used logical symbols. However, mathematical logic theory demonstrates that these symbols are entirely equivalent to appropriate combinations of existing symbols and can make many formulas more concise [37]. Therefore, they are introduced as defined symbols in subsequent discussions.
@@ Appendix B Finite Automaton States
Using the formal expression of states, we can study in depth the special and important object of finite automata [38].
@@
Proof: Let $M=\left\langle Q,R,U,\delta,\lambda\right\rangle$ be a finite automaton, and $T=\left\{t_{i}\middle|i=1,\ldots,n\right\}$ the (strictly increasing) time points at which $M$ performs input, output, or state‐transition actions. By finite‐automaton theory [38], $Q=\left\{q_{t_{i}}\middle|i=1,\ldots,n-1\right\}$, $R=\left\{r_{t_{i}}\middle|i=2,\ldots,n\right\}$, and $U=\left\{u_{t_{i}}\middle|i=1,\ldots,n\right\}$ are nonempty finite sets of input symbols, output symbols, and states, respectively;
@@
$\varphi_{\lambda}\left(t_{i}\right)=$ $\ \varphi_{U}(t_{i})\land\varphi_{I}(t_{i})\rightarrow\exists\lambda(u_{t_{i}},q_{t_{i}})\land$ (B.5)
$\mathit{Eq}^{2}(r_{t_{i+1}},\lambda(u_{t_{i}},q_{t_{i}}))\land$
$\mathit{IsElementof}^{2}(r_{t_{i+1}},R)\land\varphi_{O}(t_{i+1})$
@@
describes $M$’s inputs, outputs, and state‐transitions over $T$. Here $Cn(X)$ denotes the logical closure of the set of WFFs $X$. This completes the proof.
@@ C.1 A Simple Feedforward Neural Network Model
We choose a feedforward neural network M with an input layer containing two inputs, a hidden layer containing three neurons, and an output layer containing two outputs (Fig. C.1). Its simplicity helps explain the essence of MLT-MF in a clear and straightforward manner.
@@ C.2 Symbol Definitions
MLT-MF uses a formal system to express the operational logic of each phase throughout the model’s full lifecycle. Defining the symbols is a prerequisite for formal modeling. To this end, Tables C.1–4 provide definitions for the constants, variables, functions, and predicates used in the formal representation of model M.
@@
Table 2: Definition of Variables
@@
$\varphi_{i}^{o}$ WFFs WFFs contained in the informational ontological state of M’s initial model
$\psi_{i}^{o}$ WFFs WFFs contained in the informational carrier state of M’s initial model
$\varphi_{i}^{t}$ WFFs WFFs contained in the training informational ontological state
$\psi_{i}^{t}$ WFFs WFFs contained in the training informational carrier state
$\varphi_{i}^{u}$ WFFs WFFs contained in the informational ontological state of M’s optimized model
$\psi_{i}^{u}$ WFFs WFFs contained in the informational carrier state of M’s optimized model
$\varphi_{i}^{q}$ WFFs WFFs contained in the input informational ontological state of M
$\psi_{i}^{q}$ WFFs WFFs contained in the input informational carrier state of M
$\varphi_{i}^{r}$ WFFs WFFs contained in the output informational ontological state of M
$\psi_{i}^{r}$ WFFs WFFs contained in the output informational carrier state of M
$\varphi_{i}^{v}$ WFFs WFFs contained in the informational ontological state of M’s re-optimized model
$\psi_{i}^{v}$ WFFs WFFs contained in the informational carrier state of M’s re-optimized model
layers natural number number of layers in the model
@@
$\mathbf{W}_{k}^{(o)}$ matrix weight matrix from the $k$-th to the $(k+1)$-th layer in the initial model
$w_{ij}^{(o)k}$ real number weight from the $j$-th neuron in the $k$-th layer to the $i$-th neuron in the $(k+1)$-th layer in the initial model
$\mathbf{b}_{k}^{(o)}$ vector bias vector of neurons in the $k$-th layer in the initial model
$b_{i}^{(o)k}$ real number bias of the $i$-th neuron in the $k$-th layer in the initial model
$\mathbf{W}_{k}^{(u)}$ matrix weight matrix from the $k$-th to the $(k+1)$-th layer in the optimized model
$w_{ij}^{(u)k}$ real number weight from the $j$-th neuron in the $k$-th layer to the $i$-th neuron in the $(k+1)$-th layer in the optimized model
$\mathbf{b}_{k}^{(u)}$ vector bias vector of neurons in the $k$-th layer in the optimized model
$b_{i}^{(u)k}$ real number bias of the $i$-th neuron in the $k$-th layer in the optimized model
$\mathbf{W}_{k}^{(v)}$ matrix weight matrix from the $k$-th to the $(k+1)$-th layer in the re-optimized model
$w_{ij}^{(v)k}$ real number weight from the $j$-th neuron in the $k$-th layer to the $i$-th neuron in the $(k+1)$-th layer in the re-optimized model
$\mathbf{b}_{k}^{(v)}$ vector bias vector of neurons in the $k$-th layer in the re-optimized model
$b_{i}^{(v)k}$ real number bias of the $i$-th neuron in the $k$-th layer in the re-optimized model
$\mathbf{W}_{k}^{(w)}$ matrix weight matrix from the $k$-th to the $(k+1)$-th layer in the further re-optimized model
$w_{ij}^{(w)k}$ real number weight from the $j$-th neuron in the $k$-th layer to the $i$-th neuron in the $(k+1)$-th layer in the further re-optimized model
$\mathbf{b}_{k}^{(w)}$ vector bias vector of neurons in the $k$-th layer in the further re-optimized model
$b_{i}^{(w)k}$ real number bias of the $i$-th neuron in the $k$-th layer in the further re-optimized model
$\mathbf{X}$ vector input vector
$\mathbf{H}$ vector hidden layer vector
$\mathbf{Y}$ vector output vector
$\mathbf{Y}_{t}$ vector output reference vector
$x_{i}$ real number the $i$-th element of the input vector
$h_{i}$ real number the $i$-th element of the hidden layer vector
$y_{i}$ real number the $i$-th element of the output vector
$y_{ti}$ real number the $i$-th element of the output reference vector
@@
$\nabla_{\mathbf{W}_{1}}L$ real number gradient of the loss function with respect to $\mathbf{W}_{1}$
$\nabla_{\mathbf{b}_{1}}L$ real number gradient of the loss function with respect to $\mathbf{b}_{1}$
$\nabla_{\mathbf{W}_{2}}L$ real number gradient of the loss function with respect to $\mathbf{W}_{2}$
$\nabla_{\mathbf{b}_{2}}L$ real number gradient of the loss function with respect to $\mathbf{b}_{2}$
$\eta$ positive real number learning rate
@@
$\text{ReLU}(x)$ 1 rectified linear unit (ReLU) activation function
$\text{Identity}(x)$ 1 identity function
@@
$\text{MSE}(\mathbf{Y},\mathbf{Y}_{t})$ 2 mean squared error between vector $\mathbf{Y}$ and $\mathbf{Y}_{t}$
$\text{ReLU}^{\prime}(x)$ 1 derivative of $\text{ReLU}(x)$
$\text{Target}(\mathbf{X})$ 1 objective function for input vector $\mathbf{X}$
@@
$\text{NetworkStructure}^{2}(\text{layers},k)$ model structure has $k$ layers
$\text{LayerSize}^{2}(k,j)$ the $k$-th layer has $j$ neurons
$\text{Eq}^{2}(x,y)$ $x$ equals $y$
$\text{InputAt}^{2}(\mathbf{X},t)$ input $\mathbf{X}$ at time $t$
$\text{HiddenAt}^{2}(\mathbf{H},t)$ hidden layer value is $\mathbf{H}$ at time $t$
$\text{OutputAt}^{2}(\mathbf{Y},t)$ output $\mathbf{Y}$ at time $t$
$\text{InputAt}^{3}(\mathbf{X},\mathbf{Y}_{t},t)$ input vector $\mathbf{X}$ and label vector $\mathbf{Y}_{t}$ at time $t$
$\text{Differentiable}^{2}(f,v)$ function $f$ is differentiable with respect to $v$
$\text{BelongsTo}^{2}(x,X)$ element $x$ belongs to set $X$
@@ C.3 Formal Representations of Information at Each Stage of Model M
$Io:S_{oo}(oo,To_{h})\rightrightarrows S_{M}(M,To_{m})$ (C.13)
Here, the ontology $oo$ represents the designer’s conceptual blueprint; ${To}_{h}$ is the time at which the model is designed; and the ontological state $S_{oo}(oo,{To}_{h})$ is a set of WFFs that describe the initial operational logic of model M:
@@
$\varphi o_{1}=$ $\ \mathit{NetworkStructure}^{2}(layers,3)\land$ (C.15)
$\mathit{LayerSize}^{2}(1,2)\land\mathit{LayerSize}^{2}(2,3)\land$
$\mathit{LayerSize}^{2}(3,2)$
  • 1.
@@
$\varphi o_{2}=$ $\ \forall i\forall j\left(\mathit{BelongsTo}^{2}\left(i,\{1,2,3\}\right)\land\mathit{BelongsTo}^{2}\left(j,\{1,2\}\right)\right$. (C.16)
$\left.\rightarrow\exists v\left(\mathit{Eq}^{2}\left(w_{ij}^{(o)1},v\right)\right)\right)$
$\varphi o_{3}=$ $\ \forall i\forall j\left(\mathit{BelongsTo}^{2}\left(i,\{1,2\}\right)\land\mathit{BelongsTo}^{2}\left(j,\{1,2,3\}\right)\right$. (C.17)
$\left.\rightarrow\exists v\left(\mathit{Eq}^{2}\left(w_{ij}^{(o)2},v\right)\right)\right)$
  • 1.
@@
$\varphi o_{7}=$ $\ \forall X\forall t\left(\mathit{InputAt}^{2}(X,t)\land\mathit{BelongsTo}^{2}\left(X,\mathbb{R}^{2}\right)\rightarrow\right$. (C.21)
$\exists H\left(\mathit{BelongsTo}^{2}\left(H,\mathbb{R}^{3}\right)\land\mathit{HiddenAt}^{2}(H,t+\delta)\land\right$.
[…]
$\left.\left.\mathit{Eq}^{2}\left(h_{i},\mathrm{ReLU}\left(\sum_{j=1}^{2}w_{ij}^{(o)1}\times x_{j}+b_{i}^{(o)1}\right)\right)\right)\right)$
$\varphi o_{8}=$ $\ \forall H\forall t\left(\mathit{HiddenAt}^{2}(H,t+\delta)\land\mathit{BelongsTo}^{2}\left(H,\mathbb{R}^{3}\right)\rightarrow\right$. (C.22)
$\exists Y\left(\mathit{BelongsTo}^{2}\left(Y,\mathbb{R}^{2}\right)\land\mathit{OutputAt}^{2}(Y,t+2\delta)\land\right$.
[…]
$\left.\left.\mathit{Eq}^{2}\left(y_{i},\mathrm{Identity}\left(\sum_{j=1}^{3}w_{ij}^{(o)2}\times h_{j}+b_{i}^{(o)2}\right)\right)\right)\right)$
  • 1.
@@
$\varphi o_{9}=$ $\ \forall X\forall Y_{t}\forall t\left(\mathit{InputAt}^{3}\left(X,Y_{t},t\right)\land\mathit{BelongsTo}^{2}\left(Y_{t},\mathbb{R}^{2}\right)\land\right$. (C.23)
$\left.\varphi o_{7}\land\varphi o_{8}\land\mathit{Eq}^{2}\left(L,\mathrm{MSE}\left(Y,Y_{t}\right)\right)\land\right$.
$\left.\mathit{Eq}^{2}\left(\mathrm{MSE}\left(Y,Y_{t}\right),\frac{1}{2}\sum_{i=1}^{2}\left(y_{i}-y_{ti}\right)^{2}\right)\right)$
@@
$\quad\land\mathit{Differentiable}^{2}(MSE(Y,Y_{t}),W_{1}^{(o)})$
$\quad\land\mathit{Differentiable}^{2}(MSE(Y,Y_{t}),b_{1}^{(o)})$
$\quad\land\mathit{Differentiable}^{2}(MSE(Y,Y_{t}),W_{2}^{(o)})$
$\quad\land\mathit{Differentiable}^{2}(MSE(Y,Y_{t}),b_{2}^{(o)})$
$\quad\land\mathit{Eq}^{2}(\nabla_{W_{1}^{(o)}}L,$
[…]
$\quad\land\mathit{Eq}^{2}(\nabla_{b_{1}^{(o)}}L,$
[…]
$\quad\land\mathit{Eq}^{2}(\nabla_{W_{2}^{(o)}}L,(Y-Y_{t}){(\mathit{ReLU}(W_{1}^{(o)}X+b_{1}^{(o)}))}^{T})$
$\quad\land\mathit{Eq}^{2}(\nabla_{b_{2}^{(o)}}L,(Y-Y_{t}))$ (C.24)
  • 1.
@@
$max({To}_{h})\leq min({To}_{m})$ (C.26)
The state $S_{M}(M,{To}_{m})$ is a noisy-free enabling mapping of the state $S_{oo}\left(oo,{To}_{h}\right)$ onto M, that is,